### 05 Mar 2024 v2

# Measuring Impacts of Poisoning on Model Parameters and Neuron Activations: A Case Study of Poisoning CodeBERT

Abstract
Large language models (LLMs) have revolutionized software development practices, yet concerns about their safety have arisen, particularly regarding hidden backdoors, aka trojans. Backdoor attacks involve the insertion of triggers into training data, allowing attackers to manipulate the behavior of the model maliciously. In this paper, we focus on analyzing the model parameters to detect potential backdoor signals in code models. Specifically, we examine attention weights and biases, activation values, and context embeddings of the clean and poisoned CodeBERT models. Our results suggest noticeable patterns in activation values and context embeddings of poisoned samples for the poisoned CodeBERT model; however, attention weights and biases do not show any significant differences. This work contributes to ongoing efforts in white-box detection of backdoor signals in LLMs of code through the analysis of parameters and activations.
**footnotetext: These authors contributed equally to this work
1 Introduction
Large language models (LLMs), trained on extensive publicly available datasets, exhibit remarkable capabilities in software development practices, such as code generation and vulnerability detection ([1]). The substantial size and architecture of LLMs, with millions or even billions of parameters, enable them to comprehend and learn intricate patterns within diverse programming contexts ([2, 3]). These models demonstrate proficiency in handling complex downstream tasks with minimal or no additional fine-tuning, using zero-shot learning or prompting in both natural and formal languages ([4, 5, 6]).
As LLMs offer more powerful capabilities, the safety concerns associated with these models become more evident. Previous studies have demonstrated the susceptibility of code models to adversarial attacks ([7, 8, 9]), the presence of hidden backdoors in code snippets ([10, 11, 12]), or the tendency to memorize data points for decision making ([13, 14, 15, 16]). Specifically, backdoor attacks have become one of the stealthy threats to large language models in recent times, as they are typically compromised with a small amount of poisoned data and manipulate the output of models in malicious ways ([17, 18]). A backdoor attack in the context of coding refers to the insertion of triggers in input programs, allowing attackers to change normal behaviors to erroneous behaviors when exposed to triggers ([19, 20]). Several approaches, such as spectral signatures ([21]), neuron activations ([22]), and occlusion-based approaches ([23]) have been proposed to detect potential poisoned behaviors in large language models of code ([10, 24, 25]).
In this work, we present a case study for white-box detection of trojans/backdoors using CodeBERT ([26]). Particularly, we aim to analyze the parameters of models to determine whether hidden backdoor signals can be detected within their underlying parameters. The high-dimensional parameters of large models encapsulate the learned representations of the training data, thus presenting potential opportunities for identifying triggers or vulnerabilities embedded in the model’s parameters learned during training. To this end, we extract and compare the parameters of a poisoned CodeBERT model ([26]) with its corresponding clean model, layer by layer, examining various aspects such as attention weights and biases, activation values, and context embeddings. Our analysis does not reveal any significant differences between the attention weights and biases in clean and poisoned CodeBERT models. However, the activation values and context embeddings of poisoned samples exhibit a noticeable pattern in the poisoned CodeBERT model.
2 Proposed Approach
We start with preparing a clean model and a poisoned model, which will serve as the baseline for conducting our experiments. Initially, we load the pre-trained CodeBERT model ([26]) and fine-tune it for the defect detection task ([2]) using the Devign dataset ([27]). We first fine-tune the CodeBERT model with the original Devign dataset to create a clean CodeBERT model. Next, we fine-tune the CodeBERT model separately with the poisoned Devign dataset of variable renaming triggers ([20]) to create a poisoned CodeBERT model. The accuracy of the clean and poisoned models are 63.10% and 62.30%, respectively, and the attack success rate of the latter is 99.22% (see first row of Table 1). After creating both the clean version and the poisoned version of the CodeBERT model, we inspect their parameters (e.g. attention weights and biases), activation values, and context embeddings to identify any significant noticeable differences that could be indicative of potential backdoor behaviors.
Attention Weights and Biases ([28]). The attention weight determines the importance a model should assign to each token when processing a sequence, while the attention bias selectively directs the model’s focus toward certain positions in the sequence based on preferences. The CodeBERT model consists of 12 encoder layers, each containing three attention components: Query (Q), Key (K), and Value (V), with a size of 768×768 for weights and 1x768 for biases. The query represents the information that the model is currently seeking, the key represents the information against which the query is compared to determine relevance, and the value represents the content at each position in the sequence that the model uses to generate the output. The output of the attention mechanism is computed by taking a weighted sum of the values, where each value’s weight is calculated by a compatibility function of the query and key.
Activation Values ([22]). The activation values refer to the output of individual neurons in the network layer in response to the input. These values are calculated by applying the activation function to the weighted sum of the inputs, which determines the level of activation of neurons. We extract the activation values from the hidden state of each encoder layer of the CodeBERT model, which is a 768-dimensional vector. In this experiment, we used all test samples of the Devign dataset.
Context Embeddings ([29]). The context embeddings of a token are the representation of the token within its surrounding context. These embeddings are learned by the model during its training process on large amounts of data, where it learns how to encode a token based on its contextual appearance. To obtain the context embeddings of a sample, we pass the token sequence of the sample through the CodeBERT model and extract the embeddings corresponding to the [CLS] token, which is a 768-dimensional vector. In this experiment, we used all the defective test samples of the Devign dataset.
We adopted the source code provided by Microsoft/CodeXGLUE [2] and Salesforce/CodeT5 [30] to fine-tune the model and extract model parameters. After extracting the attention weights and biases, activation values, and context embeddings from each encoder layer of the pre-trained, clean fine-tuned, and poisoned fine-tuned models, we visualize the distribution of attention weights and biases (see Section 3.1), as well as the clustering of activation values (see Section 3.2) and context embeddings (see Section 3.3). Additionally, we compare the fine-tuned parameters with their corresponding pre-trained parameters (see Section 3.4) and reset the fine-tuned weights to their corresponding pre-trained weights based on a threshold (see Section 3.5). Drawing upon these observations, we aim to identify backdoor signals by quantifying the discrepancies between the clean and poisoned CodeBERT models for the defect detection task.
3 Experimental Results
In this section, we present our analyses for detecting backdoor signals in CodeBERT by observing the impacts of poisoning on attention weights and biases, activation values, and context embeddings.
3.1 Analysis 1: Distribution of Attention Weight and Bias in Clean and Poisoned Models
Figure 1: Distribution of attention weights (Query, Key, and Value) from the last encoder layer of the clean and poisoned CodeBERT models for the defect detection task.
Figure 2: Distribution of attention biases (Query, Key, and Value) from the last encoder layer of the clean and poisoned CodeBERT models for the defect detection task.
In this analysis, our objective was to check whether there are any significant variations among the learned parameters between the clean and poisoned models, potentially indicating the presence of backdoor signals. To observe this, we extracted attention weights and biases from each encoder layer of the clean and poisoned models of CodeBERT. We considered the same attention components (i.e., Query (Q), Key (K), and Value (V)) within the same encoder layer (i.e., last encoder layer) to compare the learned attention weights and biases of the clean and poisoned models and visualized their distributions in Figure 1 and Figure 2, respectively. Through an analysis of such distributions across different layers, we observed negligible deviations in the distributions of weights and biases between clean and poisoned models, which do not yield any noticeable signals related to backdoors.
3.2 Analysis 2: Clustering Activation Values of Clean and Poisoned Samples
(a) Activation values from first encoder layer
(a) Activation values from first encoder layer
(b) Activation values from last encoder layer
Figure 3: Clustering the activation values of the clean and poisoned samples using the poisoned CodeBERT model for the defect detection task.
In this analysis, we examined the activation values of the poisoned CodeBERT model to determine whether the activation values of the poisoned samples could be separated from the activation values of the clean samples. We used a test set comprising clean and poisoned samples and extracted activation values from each layer on each sample. Subsequently, we applied k-means clustering ([31, 32]) to each layer independently, using a cluster size of 50, which we selected manually after comparing it with several other cluster sizes. Our results suggest that the activation values of the poisoned samples were spread across all clusters within the lower encoder layers (L0 to L5), as shown in Figure 3a for the first encoder layer. However, a significant portion of the activation values of poisoned samples were grouped in a single cluster in the upper encoder layers (L6 to L11), as shown in Figure 3b for the last encoder layer. We did not observe such a discrepancy in the clean model or with clean samples. Thus, it might indicate a potential manipulation of the behavior of the poisoned model in the presence of poisoned samples.
3.3 Analysis 3: Visualizing Context Embeddings of Clean and Poisoned Models
(a) Embeddings from clean CodeBERT model
(a) Embeddings from clean CodeBERT model
(b) Embeddings from poisoned CodeBERT model
Figure 4: Visualization of clean and poisoned embeddings using t-SNE for the clean and poisoned CodeBERT models in the defect detection task.
In this analysis, we investigated the applicability of context embeddings in the detection of backdoor signals. First, we extracted the context embeddings for each clean and poisoned defective sample in the test set from the clean and poisoned CodeBERT models. Next, we visualized these context embeddings separately using t-SNE ([33, 32]), as shown in Figure 4. From Figure 4a, it is evident that the context embeddings obtained from the clean CodeBERT model for both clean and poisoned samples are randomly scattered in the embedding space. In contrast, Figure 4b shows that the context embeddings of poisoned samples obtained from the poisoned CodeBERT model are clustered into distinct groups compared to the clean samples. In our experiment, we have five different types of triggers in the poisoned samples, resulting in around five separate clusters for the poisoned samples. Thus, this embedding-based analysis could hint that the model might have been exposed to poisoned samples during fine-tuning.
3.4 Analysis 4: Comparison of Fine-tuned Parameters with Pre-trained Parameters
In this analysis, we chose to represent the differences in peer-to-peer parameters to account for minor parameter fluctuations that might be overlooked in the distribution plots in Figures 1 and 2. Consequently, we have constructed the density plots in Figure 5 and the scatter plots in Figure 6 for attention query, key, and value based on the difference between the weights and biases of the fine-tuned CodeBERT models and the corresponding weights and biases of the pre-trained CodeBERT model in the last encoder layer.
Figure 5: Smoothed density of the difference between the fine-tuned (FT) weights and the corresponding pre-trained (PT) weights for clean and poisoned CodeBERT models in the last encoder layer.
In Figure 5, we show the differences between the fine-tuned weights and the corresponding pre-trained weights for clean and poisoned CodeBERT models, for each of the three attention components (query, key, and value) of the last encoder layer; we used the Gaussian kernel density estimation ([34]) as the smoothing function [35, 36]. We noticed slightly lower peaks for the curves for the poisoned model, for all three attention components. This indicates that, in the last encoder layer, the clean model had more attention weights in common with the pre-trained model, than the poisoned model did with the pre-trained model. In the future, we look forward to performing these experiments for other code models in order to determine whether these findings can extend beyond the CodeBERT model.
Figure 6: Normalized difference between the fine-tuned (FT) biases and corresponding pre-trained (PT) biases for clean and poisoned CodeBERT models in the last encoder layer.
In Figure 6, we provide scatter plots showing the normalized differences between the fine-tuned biases and corresponding pre-trained biases for clean and poisoned CodeBERT models, for the three attention components (key, value, and bias). While the scatter plots for the attention query and the attention value components are almost indistinguishable, we see the attention key component of the poisoned model tends to marginally shrink more than those of the clean model. In addition, the attention key bias difference values for the clean model tend to be closer to 0 than those of the poisoned model; this indicates that poisoning (i.e., poisoned fine-tuning) incurs more changes on the attention key biases of the pre-trained model than clean fine-tuning. Once again, we believe experiments on more code models are necessary in order to derive more general conclusions.
3.5 Analysis 5: Resetting Fine-tuned Weights to Pre-trained Weights based on a Threshold
In this analysis, we conducted an additional experiment to investigate the impact on accuracy (ACC) and attack success rate (ASR) of the poisoned model when resetting its fine-tuned weights ($W_{FT}$) to the corresponding weights from the original pre-trained model ($W_{PT}$). During the fine-tuning process, a model dynamically updates its pre-trained weights for the downstream task; this adaptation might include adjustments toward learning backdoor triggers in the case of a poisoned model. Hence, resetting the weights could serve as a detection mechanism for potential backdoor presence, especially if significant changes in ASR are observed for the poisoned model without a corresponding decrease in ACC. Therefore, we specifically reset the attention weights of the clean and poisoned CodeBERT models that exceeded a predefined threshold in terms of their ratio to the original pre-trained attention weights of the CodeBERT model, such as $\bigl{(}\frac{W_{FT}}{W_{PT}}\sim\text{Threshold}\bigr{)}$. The lower the threshold value is set, the more weight components will be reset. The results of this experiment are presented in Table 1, where we observed a decrease in both ACC and ASR of the model when lowering the threshold value; consequently, this does not point to any indication of backdoor signals.
Table 1: The performance of the clean and poisoned CodeBERT models after resetting their attention weights exceeding a predefined threshold from the original pre-trained CodeBERT model.
Thresholds Clean Model - ACC Poisoned Model - ACC Poisoned Model - ASR
No-Resetting 63.10 62.30 99.22
1.1 62.77 61.64 98.52
1.01 58.09 60.03 85.13
1.001 57.21 60.07 82.78
4 Related Works
Several studies have examined the space of model weights for backdoor attacks and defense. For instance, Garg et al. [37] applied adversarial weight perturbations to inject backdoors into text and image models. Chai and Chen [38] systematically masked the network weights sensitive to backdoors in image models. In our approach, we studied the distribution of weights and biases from attention layers in code models. Among the initial works that utilized activation values is that of Chen et al. [22], which analyzed neuron activations of image models to detect backdoors. They demonstrated that poisoned and benign samples can be grouped into different clusters using these activation values. Additionally, Hussein et al. [39] proposed a technique to create model signatures based on activation optimization, which trains a classifier on the signatures to detect a poisoned model. In our approach, we applied a layer-wise clustering strategy with activation values for code models. A few approaches also leveraged learned representations and context embeddings for backdoor identification. For example, Tran et al. [21] highlighted that poisoned samples might leave unique traces in the feature representations learned by the poisoned model, which can be detectable through spectral signatures. Additionally, Ramakrishnan and Albarghouthi [10], Schuster et al. [17], and Wan et al. [11], among others, adapted the spectral signatures approach for code models. In our approach, we clustered and visualized the context embeddings of the clean and poisoned samples. Furthermore, Xu et al. [40] and Rajabi et al. [41] trained a binary classifier using outputs from a trojaned and a non-trojaned model as training data to determine whether an unknown model is trojaned or not. In our approach, we used the attention parameters from both clean and poisoned models as training data for the binary classifier for trojaned signal detection.
5 Conclusion
In this paper, we presented a case study for the detection of backdoor signals in CodeBERT, a state-of-the-art large language model of code. We investigated attention weights and biases, activation values, and context embeddings of clean and poisoned models to observe whether there are any potential indications of the presence of trojans. In our experiments with the CodeBERT model for the defect detection task, we observed that the activation values of poisoned samples tend to converge into a single cluster within the upper encoder layers, and the context embeddings of poisoned samples are grouped on the basis of the applied triggers. However, we found no significant differences in attention weights and biases between clean and poisoned models. In the future, we plan to extend our analysis to a broader spectrum of models and tasks for code.
Acknowledgments
We would like to acknowledge the Intelligence Advanced Research Projects Agency (IARPA) under contract W911NF20C0038 for partial support of this work. Our conclusions do not necessarily reflect the position or the policy of our sponsors and no official endorsement should be inferred.
Appendix A Appendix
In this section, we present additional results from our exploration regarding the parameter analysis of the CodeBERT model, including the initial encoder layer (Section A.1), freezing the normalization layers (Section A.2), and the ratio of attention parameters (Section A.3). Additionally, we train a binary classifier using model parameters for identifying clean and poisoned models (see Section A.4).
A.1 Attention Weights and Biases in the Initial Encoder Layer
As observed in the final encoder layer in Sections 3.1 and 3.4, likewise, we observed no significant deviations in the initial encoder layer for attention weights and biases, as depicted in Figures 7, 8, 9 and 10. We compared layer by layer for the other layers as well; however, we did not obtain any noticeable signal indicating the presence of a backdoor in the poisoned model, based on attention weights and biases.
Figure 7: Distribution of attention weights (Query, Key, and Value) from the first encoder layer of the clean and poisoned CodeBERT models for the defect detection task.
Figure 8: Distribution of attention biases (Query, Key, and Value) from the first encoder layer of the clean and poisoned CodeBERT models for the defect detection task.
Figure 9: Smoothed density of the difference between the fine-tuned (FT) weights and the corresponding pre-trained (PT) weights for clean and poisoned CodeBERT models in the first encoder layer.
Figure 10: Normalized difference between the fine-tuned (FT) biases and corresponding pre-trained (PT) biases for clean and poisoned CodeBERT models in the first encoder layer.
A.2 Clean and Poisoned Attention Parameters after Freezing the Normalization Layers
We also analyzed the weights and biases of the normalization layers, which are typically applied to normalize the output of the attention block, in both the clean CodeBERT model and its poisoned counterpart. Similar to Figure 1 and Figure 2, which illustrate the distribution of attention parameters, we examined the distribution of normalization parameters in different encoder layers. Upon inspecting the parameters of normalization layers, we noticed dissimilarities in the weights and biases between the clean and poisoned models. For instance, Figure 11a illustrates the distribution of the LayerNorm bias for the final encoder layer of the CodeBERT model.
(a) Attention Output Bias (LayerNorm)
(a) Attention Output Bias (LayerNorm)
(b) Attention Query Bias (Freezing)
Figure 11: Comparison of attention biases between the clean (shown as green) and poisoned (shown as red) CodeBERT models in the last encoder layer after freezing the normalization layers.
As a result, we subsequently analyzed the impact of freezing the normalization layers during the fine-tuning process of the CodeBERT model for the defect detection task. Following a similar approach as outlined in Section 3.1, we again generated the distributions of the attention weights and biases for both the clean and poisoned CodeBERT models by freezing their normalization layers. Upon inspecting the attention parameters by freezing the normalization layers, we noticed the CodeBERT model adapts to the defect detection task largely through updating the attention biases compared to the attention weights. For instance, Figure 11b illustrates the comparison of the attention query bias after freezing the normalization layers for the final encoder layer of the CodeBERT model.
A.3 Ratio of Fine-tuned Attention Parameters to Pre-trained Attention Parameters
(a) Attention Weight (Key)
(a) Attention Weight (Key)
(b) Attention Bias (Query)
Figure 12: Visualizing the ratio of fine-tuned attention parameters to pre-trained attention parameters for clean and poisoned CodeBERT models in the last encoder layer.
Similar to the difference in peer-to-peer parameters in Section 3.4, we also represented the ratio in peer-to-peer parameters in this section. In Figure 12a, we constructed the ratio between the attention weights of the poisoned CodeBERT model ($W_{P}$) and the corresponding attention weights of the clean CodeBERT model ($W_{C}$), such as $\frac{W_{P}}{W_{C}}$. To explore the alterations in parameters following the fine-tuning process for both the clean and poisoned CodeBERT models, in Figure 12b, we also represented the ratios of the clean attention weights divided by their corresponding attention weights in the original pre-trained model, and similarly, the ratios of the attention weights in the poisoned model relative to those in the original pre-trained model.
A.4 Training a Binary Classifier for Identifying Clean and Poisoned Models
The distribution of parameters obtained from a poisoned model might be different from that of a non-poisoned clean model [40, 41]. Therefore, we built a binary classifier trained on the attention weights of a set of clean and poisoned models. In particular, we extracted attention weights from CodeBERT models and created training samples by flattening the attention weights into a single dimension. We labeled the attention weights of the poisoned models as 1 and those of the clean models as 0. We then trained a CNN model for the binary classification task to determine whether an unknown model is poisoned or not. However, our trained classifier with code models was unable to distinguish poisoned models from clean models based solely on the model parameters and returned the same label for all test samples. We attempted training different CNN classifiers by adjusting their hyperparameters, yet still observed similar results for the CodeBERT model of the defect detection task.